OpenStudy (anonymous):

solve the following equations for all possible values of x: 1) 5sinx+1=3sinx 2) 7cosx + 9= -2cosx 3) cot (3x/2) = -√3

OpenStudy (anonymous):

please answer at least one! [:

OpenStudy (mertsj):

\[2sinx=-1\] \[sinx=\frac{-1}{2}\] \[x=\frac{7\pi}{6}, \frac{11\pi}{6} + 2\pi n\]

OpenStudy (mertsj):

\[9\cos x=-9\] \[\cos x=-1\] \[x=\pi +2\pi n\]

OpenStudy (mertsj):

\[\tan \frac{3x}{2}=-\frac{\sqrt{3}}{3}\] \[\frac{3x}{2}=\frac{5\pi}{6}, \frac{11\pi}{6}, \frac{17\pi}{6}\] \[x=\frac{10\pi}{18},\frac{22\pi}{18}, \frac{34\pi}{18} + 2\pi n\]